OpenStudy (anonymous):

Three 1.0pF capacitors are charged separately to the potential difference of 100,200 and 300 volts. The capacitors are then joined in parallels.What is the resultant potential difference.

OpenStudy (nikvist):

\[C_1=C_2=C_3=C=1\,pF\quad,\quad U_1=100V\enspace,\enspace U_2=200V\enspace,\enspace U_3=300V\]\[W_0=\frac{1}{2}C\cdot\left(U_1^2+U_2^2+U_3^2\right)\quad,\quad W=\frac{1}{2}C\cdot 3U^2\]\[W=W_0\quad\Rightarrow\quad 3U^2=U_1^2+U_2^2+U_3^2\]\[U=\sqrt{\frac{U_1^2+U_2^2+U_3^2}{3}}=216.025\,V\]

OpenStudy (anonymous):

hey nikvist how do add subscript and superscript in your posts.

OpenStudy (nikvist):

\[\mbox{X_2}\quad\Rightarrow\quad X_2\]\[\mbox{X^2}\quad\Rightarrow\quad X^2\]

OpenStudy (anonymous):

A_4 A^4

OpenStudy (anonymous):

not working?

OpenStudy (anonymous):

yea got it

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

\[Thanq^6_3\]